OpenStudy (anonymous):

a package of supplies is to be dropped from an airplane so that it hits the ground at a designated spot near some campers. The airplane, moving horizontally at a constant velocity of 140 km/hr, approaches the spot at an altitude of 0.5 km above level ground. Having the designated point in sight, the pilot prepares to drop the package. How far horizontally from the drop point should the pilot release the package? What is the location of the plane when the package hits the ground?

OpenStudy (anonymous):

3.85km (approx)

OpenStudy (anonymous):

when the package hits the ground, the plane will be vertically above the package

OpenStudy (anonymous):

how do you figure this out though. what are the steps?

OpenStudy (anonymous):

from kinematics u can find (using y axis eqns) the total time of fall. let it be t2. and since no force acts in the horiz. direction, so dist.=speed * time( which's t2 in our case). thus req. horiz. dist = 140km/hr*(t2 sec).

OpenStudy (anonymous):

For Vertical component: Using \[s=ut +0.5at^2\] \[0.5(1000)=0(t)+0.5(9.81)t^2\] \[t=\sqrt{1000/9.81}\] For Horizontal component: Using \[s =vt\] \[s=[140(1000)]/[60(60)]\times \sqrt{1000/9.81}\]\[=392 m\]

OpenStudy (anonymous):

thank you so much!